OpenStudy (anonymous):

Write the equation of the parabola in vertex form vertex(-4,3) point(4,131. a. y=2(x+4)^2+3 b.y=2(x-4)^2+3 c.y=4(x-4)^2+3 d.y=131(x+4)^2-3

OpenStudy (anonymous):

Is the given point (4, 131)?

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

As the vertex is (-4, 3), let the parabola is (y-3)=a(x+4)^2. As the point (4, 131) should lie on the parabola, the value of 'a' must be chosen such that the equation is satisfied with x=4 and y=131 are substituted. So, we should have (131-3)=a(4+4)^2. This gives 128=64a. Hence 'a' should be 2. The first option is correct. Hope this helps

OpenStudy (anonymous):

Thank you it cleared it up

OpenStudy (anonymous):

wait is it A then